# A Longitudinal Study on the Attitudes of Gay Men in Beijing Towards Gay Social Media Platforms: Lonely Souls in the Digital Concrete Jungle

### v2

@@ Abstract.
Over the past decade, specialized social networking applications have become a cornerstone of life for many gay men in China. This paper employs a longitudinal mixed-methods approach to investigate how Chinese men who have sex with men (MSM) have shifted their attitudes toward these platforms between approximately 2013 and 2023. Drawing on archival analysis of online discourses, a quantitative survey of 412 participants, and in-depth semi-structured interviews with 32 participants, we trace the complex trajectory of this evolution. Our findings reveal a clear pattern: from the initial embrace of these applications as revolutionary tools for community building and identity affirmation (2014–2017), to a period of growing ambivalence and critique centered on commercialization, “hookup culture,” and multiple forms of discrimination (2017–2020), and finally to the present era (2020–2023), characterized by pragmatic, fragmented, yet simultaneously critical and reconstructive uses. Today, users strategically employ a repertoire of applications—including global platforms (e.g., Grindr and Tinder), domestic mainstream platforms (e.g., Blued), and niche alternatives (e.g., Aloha)—to fulfill differentiated needs. We develop a detailed temporal framework to capture this attitudinal evolution and discuss its design implications for creating more supportive, secure, and community-oriented digital environments for marginalized groups.
@@
$c$s: Human-centered computing Social networking sitesccs: Collaborative and social computingccs: Collaborative and social computing systems and toolsccs: Social networking sites
@@ 1. Introduction
This research examines how Chinese gay men’s attitudes toward such platforms have evolved over the past decade. Specifically, it addresses the following questions:
@@
The study provides detailed empirical evidence for understanding how marginalized groups adapt to technology under conditions of high sociopolitical pressure. It contributes a unique longitudinal perspective to the field of Human–Computer Interaction (HCI) and offers valuable implications for designing more resilient and secure digital spaces.
2. Related Work: Situating a Decade of User Experience in HCI
The decade-long experience of Chinese gay men(li2024online; wang2022gay; cui2022we; o2025finding; winchester2012use; cui2022so; bardzell2020join; xie2025futures; li2023we; ma2024evaluating; bardzell2011pleasure; kong2018gay) with social apps, as documented in this study, provides rich and unique empirical evidence for several core research areas in Human-Computer Interaction (HCI) and Computer-Supported Cooperative Work (CSCW). The trajectory of user attitudes—from initial feelings of liberation to later critical reflection and current strategic use—not only corroborates existing theories but also offers new insights and challenges their applicability within marginalized communities and high-pressure political contexts. This section situates our findings within key theoretical frameworks in HCI to articulate the study’s theoretical contributions to the field.
@@ 2.2. The Shaping Role of Platform Affordances in Queer Sociality
  • Brad(myers1994challenges)’s grid interface afforded immediate encounters based on geographical proximity, greatly facilitating users seeking offline contact.
  • Swipe-based matching mechanism(li2020swipe) afforded a connection model based on mutual selection, adding a layer of deliberation and consent to interactions.
@@
These different design choices provided users with a variety of “digital homes”(desjardins2015investigating) in the early stages of the study, meeting their diverse needs from finding sexual partners to building friendships and communities.
Similarly, platform design can also afford negative social interactions(lakey1994negative; rook1984negative). The discrimination and harassment phenomena widely reported by users in the middle to late stages of this study (2017-2020) are also closely related to platform design choices. For example, many platforms set fixed role tags (e.g., “1/0”) in profiles, which, while convenient for quick filtering, also solidified and reinforced stereotypes and role discrimination within the community. Recommendation algorithms that overemphasize singular dimensions like appearance can afford body shaming and anxiety(schulenberg2023creepy). And while platforms offer convenient anonymity and instant messaging, a lack of effective identity verification and community governance mechanisms can afford deceptive behavior and online harassment.
A core finding of this study—the continuous rise in user satisfaction with app usability (e.g., simple interface, reasonable functions) alongside a sharp decline in satisfaction with sociability outcomes (e.g., sense of belonging, long-term relationships)—can be profoundly explained by affordance theory. This phenomenon reveals a disconnect(nguyen2021managing) between usability affordances and sociability affordances in platform design. As platforms accelerated their commercialization, their design iterations increasingly focused on optimizing user retention and monetization, such as through algorithmic recommendations and live-streaming gifts. These improvements may have enhanced the operational fluency of certain interfaces (usability affordances) but did so at the expense of the community’s healthy ecosystem, weakening the sociability affordances that promote deep connections and community trust. The results strongly demonstrate that for social platforms serving marginalized communities, technological ”progress” that contradicts the core social needs of the community will ultimately lead to user alienation and disillusionment.
@@ 2.3. Extending Social Transition Machinery to Longitudinal Identity Management
The user behavior pattern observed in the later stage of this study (2020-2023)—strategically combining multiple domestic and international apps to meet different needs—is highly consistent with the theory of ”social transition machinery” proposed by HCI scholar Oliver Haimson(haimson2018social). Furthermore, the longitudinal data from this study provide a unique perspective for extending this theory.
Haimson’s framework suggests that individuals undergoing major life transitions (such as a queer gender transition) utilize multiple social media platforms with different functions and separate audiences, combining them into a coordinated “machine” to manage the complex process of identity reconstruction. For example, one might use one platform for private identity(10.1016/j.intcom.2011.02.002) exploration and community support, while using another for public announcements to a broader social network.
@@
In summary, this study provides the HCI and CSCW fields with a profound case study of the technology acceptance lifecycle for a marginalized community. It warns us that for designing for marginalized communities, merely providing the affordance of ”connection” is far from sufficient. If a design fails to actively and continuously commit to building an infrastructure of care—one that can resist commercial erosion, promote community health, and empower users to manage their own safety and well-being—then the initial promise of empowerment may eventually evolve into new forms of harm and exploitation. Digital platforms could then become venues that replicate and even amplify offline social injustices, rather than tools of liberation. This lays a solid theoretical and empirical foundation for the subsequent discussion of design implications with socio-cultural sensitivity.
@@ 3. Method
Over a period of nine years, we employed a mixed-methods approach, combining quantitative(adams2008qualititative) and qualitative(mcdonald2019reliability) research, to investigate the changing attitudes of 17 Chinese gay men towards gay social networking apps. The quantitative component utilized a usability scale to quantify participants’ assessments of the basic viability of these apps. The qualitative research consisted of semi-structured interviews covering: 1) basic user information (age, education, urban/rural residence, coming-out status); 2) types, frequency, and purposes of gay social media usage; 3) user experience with gay social apps, including functional design, interactive atmosphere, privacy and security, mental health, and social identity; and 4) open-ended questions about the limitations and potential development of existing apps, as well as expectations for digital technology, the future of the gay community, and their personal futures.
@@ 3.1. Participants
Ethics. The study protocol was reviewed and approved by the Institutional Review Board (IRB) of an anonymized university. All participants were fully informed of the research objectives, procedures, and potential impacts. They were reminded of their rights to withdraw at any time without justification, to request deletion of their data, and to review any information collected about them. All data were anonymized, and pseudonymous participant IDs were assigned. Interviewers underwent extensive training in qualitative methods and LGBTQ+ cultural competence to ensure respectful engagement. Each participant received 30 RMB (approximately \$4.50 USD) for each completed interview session.
@@ 3.2. Procedure
We conducted four rounds of data collection in April 2014, 2017, 2020, and 2023. Each round consisted of two components:
@@ 3.4. Data Analysis
Quantitative data were analyzed using descriptive statistics and exploratory factor analysis (EFA) in SPSS to identify attitude clusters. Qualitative data were transcribed and analyzed using thematic analysis. Coding followed an iterative process, starting with open coding and then grouping by themes.
@@ 4. Results
Table 3. Participant Information Across Four Waves of Interviews
ID Age (2014) 2014 2017 2020 2023
1 19 High school Student U * * * * * * Bachelor Official *
3 22 Bachelor Student U * * * Bachelor Programmer R * * *
5 24 Bachelor Chef U Bachelor Chef U Bachelor Chef U Bachelor Chef U
7 27 Bachelor Labor U Bachelor Labor U Bachelor Labor U Bachelor Labor U
12 40 Bachelor Teacher U Bachelor Teacher U Bachelor Teacher U * * *
13 37 Master Student U PhD Teacher U PhD Unemployed U PhD Unemployed U
14 39 PhD Teacher U PhD Teacher U PhD Teacher U PhD Unemployed U
16 24 Bachelor Waiter U Bachelor Waiter U Bachelor Cleaning U * * *
17 24 High school Deliveryman R High school U * * * * * *
@@ 4.1. Quantitative Results
Qualitative Results
The qualitative findings largely corroborate the quantitative results. Initially, participants viewed gay social networking apps (GSN apps) as revolutionary tools for community building and identity affirmation. As a result, the apps were warmly received (2014–2017). From 2017 to 2020, technical architectures continued to improve. Basic usability also improved. Yet participants increasingly expressed disapproval on multiple fronts: commercialization; “fast-food” interactions; hookup culture; lookism; role-based discrimination; and pervasive deceptive practices. Reliance on GSN apps began to decline during this period.
In the 2020–2023 interviews, disappointment became salient. Participants displayed clearer segmentation and a pragmatic stance. Many adopted international GSN apps (e.g., Grindr, Tinder). Use of domestic mainstream apps (e.g., Blued, Aloha) decreased.

### v3

@@ Abstract.
Over the past decade, specialized social networking applications have become a cornerstone of life for many gay men in major metropolitan areas in China, such as Beijing. This paper employs a longitudinal mixed-methods approach to investigate how men in Beijing who have sex with men (MSM) have shifted their attitudes toward these platforms between approximately 2014 and 2023. Drawing on archival analysis of online discourses and in-depth semi-structured interviews with 17 participants, we trace the complex trajectory of this evolution. Our findings reveal a clear pattern: from the initial embrace of these applications as tools for community building and identity affirmation (2014–2017), to a period of growing ambivalence and critique centered on commercialization, “hookup culture,” and forms of discrimination (2017–2020), and finally to the present era (2020–2023), characterized by pragmatic, fragmented, yet simultaneously critical and reconstructive uses. Today, users strategically employ a repertoire of applications—including global platforms (e.g., Grindr and Tinder), domestic mainstream platforms (e.g., Blued), and niche alternatives (e.g., Aloha)—to fulfill differentiated needs. We develop a detailed temporal framework theorizing from this urban context to capture this attitudinal evolution and to discuss design implications for creating more supportive, secure, and community-oriented digital environments for sexual minority users in China. While our findings are grounded in a Beijing-based longitudinal panel, Beijing represents an important urban environment where many digital practices among Chinese MSM have been pioneered and observed. Our theoretical claims should therefore be understood as emerging from this specific socio-urban context.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Social networking sitesccs: Collaborative and social computingccs: Collaborative and social computing systems and toolsccs: Social networking sites
@@ 1. Introduction
We adopt portfolio strategy as our primary explanatory lens: Chinese gay men continuously reconfigure combinations of apps in response to persistent sociocultural pressures, with platform affordances serving as contextual triggers rather than sole drivers of change (haimson2018social). This research examines how Chinese gay men’s attitudes toward such platforms have evolved over the past decade. Specifically, it addresses the following questions:
@@
By grounding a temporal account of queer digital adaptation in established theories of performative identity, privacy, and affordances, this study contributes a theoretically integrated understanding of how marginalized users negotiate sociotechnical constraints across time. The study provides detailed empirical evidence for understanding how marginalized groups adapt to technology under conditions of high sociopolitical pressure. It contributes a unique longitudinal perspective to the field of Human–Computer Interaction (HCI) and offers valuable implications for designing more resilient and secure digital spaces.
2. Related Work: LGBTQ+ Identities, Sociotechnical Risks, and Platformized Sociality
The decade-long experience of Chinese gay men with social apps (kong2018gay; bardzell2011pleasure; ma2024evaluating; li2023we; xie2025futures), as documented in this study, provides rich and unique empirical evidence for several core research areas in Human-Computer Interaction (HCI) (cui2022we; o2025finding; winchester2012use; cui2022so) and Computer-Supported Cooperative Work (CSCW) (bardzell2020join; wang2022gay). The trajectory of user attitudes—from initial feelings of liberation to later critical reflection and current strategic use—not only corroborates existing theories but also offers new insights and challenges their applicability within marginalized communities and high-pressure political contexts. This section situates our findings within key theoretical frameworks in HCI to articulate the study’s theoretical contributions to the field. Early work on online identity highlights tensions between authenticity, privacy, and impression management in digital sociality (whitty2008revealing). Building on this foundation, LGBTQ+ HCI scholarship demonstrates that platforms are central sites for identity work, disclosure, and visibility negotiation (blackwell2016lgbt; devito2018too; 10.1145/3274313; 10.1145/3432951; hardy2022lgbtq; 10.1145/3462204.3481786; 10.1145/3613904.3642482; 10.1145/3643834.3661564). Social support emerges across diverse sociotechnical environments—from fandom communities (10.1145/3359256) to VR spaces (10.1145/3411763.3451673; 10.1145/3544548.3581530), information activism in rural contexts (jonas2024better; 10.1145/3686991), and identity-based governance in online communities (10.1145/3710900; 10.1145/3406522.3446040). Yet LGBTQ+ participation online is shaped by moderation politics and algorithmic curation biases (10.1145/3479610; 10.1145/3613904.3641949; 10.1145/3613904.3642509; 10.1145/3706598.3713618). These systems produce “algorithmic exclusion” (10.1145/3432951), particularly for trans and disabled creators (10.1145/3613904.3641949), and reinforce homonormative aesthetics in curated feeds (10.1145/3706598.3713618). Queer joy flourishes despite these harms (10.1145/3706598.3713592), reflecting HCI calls to move beyond deficit framings (10.1145/3673229). Research increasingly critiques commercial infrastructures that marginalize queer self-expression and desires for solidarity (10.1145/3706598.3713136). Affordances of LGBTQ-specific and mainstream social platforms shape relational and dating practices (10.1145/3229434.3229460; yeo2016relationships; 10.1145/3613904.3642482). Even non-dating platforms enable romantic seeking and community building (10.1145/3555189). HCI scholarship conceptualizes these dynamics as cross-platform “personal social media ecosystems” (10.1145/3274313), where LGBTQ+ users tactically modulate visibility and boundary management across contexts (10.1145/3449162; blank2014new). China-based research further highlights how queer users negotiate stigmatization and censorship in platformized ecosystems (10.1145/3491102.3517624; 10.1145/3555189; 10.1145/3706598.3714013). Mainstream and LGBTQ-specific apps alike require strategic self-presentation to mitigate outing and harassment risks (10.1145/2930971.2930973). These sociotechnical constraints shape both dating and identity development for Chinese queer communities. Taken together, this literature foregrounds a persistent tension: platforms can simultaneously afford connection, vulnerability, and sociotechnical adaptation. However, existing research is primarily cross-sectional and tends to focus on single platforms. Very little is known about how queer users’ attitudes and strategies evolve longitudinally under shifting sociocultural and platform conditions. We address this gap through a multi-wave study of Beijing gay men’s changing perceptions of gay social media platforms.
@@ 2.1. Managing Identity in Networked Publics: From Self-Presentation to Strategic Visibility
Prior research on online dating has shown how dramaturgical self-presentation shapes romantic and sexual relationship development in networked environments (whitty2008revealing), and how users often navigate a “privacy paradox” in which desires for connection coexist with concerns about surveillance and disclosure (blank2014new). These insights help contextualize our participants’ shifting negotiations between visibility, intimacy, and safety over time.
@@ 2.2. The Shaping Role of Platform Affordances in Queer Sociality
  • Myers (myers1994challenges)’s grid interface afforded immediate encounters based on geographical proximity, greatly facilitating users seeking offline contact.
  • Swipe-based matching mechanism (li2020swipe) afforded a connection model based on mutual selection, adding a layer of deliberation and consent to interactions.[7]
@@
These different design choices provided users with a variety of “digital homes” (desjardins2015investigating) in the early stages of the study, meeting their diverse needs from finding sexual partners to building friendships and communities. Similarly, platform design can also afford negative social interactions (lakey1994negative; rook1984negative). The discrimination and harassment phenomena widely reported by users in the middle to late stages of this study (2017-2020) are also closely related to platform design choices. For example, many platforms set fixed role tags (e.g., “1/0”) in profiles, which, while convenient for quick filtering, also solidified and reinforced stereotypes and role discrimination within the community. Recommendation algorithms that overemphasize singular dimensions like appearance can afford body shaming and anxiety (schulenberg2023creepy). And while platforms offer convenient anonymity and instant messaging, a lack of effective identity verification and community governance mechanisms can afford deceptive behavior and online harassment.
A core finding of this study—the continuous rise in user satisfaction with app usability (e.g., simple interface, reasonable functions) alongside a sharp decline in satisfaction with sociability outcomes (e.g., sense of belonging, long-term relationships)—can be profoundly explained by affordance theory. This phenomenon reveals a disconnect (nguyen2021managing) between usability affordances and sociability affordances in platform design. As platforms accelerated their commercialization, their design iterations increasingly focused on optimizing user retention and monetization, such as through algorithmic recommendations and live-streaming gifts. These improvements may have enhanced the operational fluency of certain interfaces (usability affordances) but did so at the expense of the community’s healthy ecosystem, weakening the sociability affordances that promote deep connections and community trust. The results strongly demonstrate that for social platforms serving marginalized communities, technological “progress” that contradicts the core social needs of the community will ultimately lead to user alienation and disillusionment. Rather than fully determining user trajectories, these affordances act as situational triggers that interact with sociocultural pressures to shape evolving strategies of use.
@@ 2.3. Extending Social Transition Machinery to Longitudinal Identity Management
The user behavior pattern observed in the later stage of this study (2020-2023)—strategically combining multiple domestic and international apps to meet different needs—is highly consistent with the theory of “social transition machinery” proposed by HCI scholar Oliver Haimson (haimson2018social). Furthermore, the longitudinal data from this study provide a unique perspective for extending this theory.
Haimson’s framework suggests that individuals undergoing major life transitions (such as a queer gender transition) utilize multiple social media platforms with different functions and separate audiences, combining them into a coordinated “machine” to manage the complex process of identity reconstruction. For example, one might use one platform for private identity (light2011hci) exploration and community support, while using another for public announcements to a broader social network.
@@
In summary, this study provides the HCI and CSCW fields with a profound case study of the technology acceptance lifecycle for a marginalized community. It warns us that for designing for marginalized communities, merely providing the affordance of “connection” is far from sufficient. If a design fails to actively and continuously commit to building an infrastructure of care—one that can resist commercial erosion, promote community health, and empower users to manage their own safety and well-being—then the initial promise of empowerment may eventually evolve into new forms of harm and exploitation. Digital platforms could then become venues that replicate and even amplify offline social injustices, rather than tools of liberation. This lays a solid theoretical and empirical foundation for the subsequent discussion of design implications with socio-cultural sensitivity.
@@ 3. Method
Over a period of nine years, we employed a mixed-methods approach, combining quantitative (adams2008qualititative) and qualitative (mcdonald2019reliability; macleod2019construction) research, to investigate the changing attitudes of 17 Chinese gay men towards gay social networking apps. The quantitative component utilized a usability scale to quantify participants’ assessments of the basic viability of these apps. The qualitative research consisted of semi-structured interviews covering: 1) basic user information (age, education, urban/rural residence, coming-out status); 2) types, frequency, and purposes of gay social media usage; 3) user experience with gay social apps, including functional design, interactive atmosphere, privacy and security, mental health, and social identity; and 4) open-ended questions about the limitations and potential development of existing apps, as well as expectations for digital technology, the future of the gay community, and their personal futures.
@@ 3.1. Participants
Table 1. Participant Information Across Four Waves of Interviews
ID Age (2014) 2014 2017 2020 2023
1 19 High school Student U * * * * * * Bachelor Official *
3 22 Bachelor Student U * * * Bachelor Programmer R * * *
5 24 Bachelor Chef U Bachelor Chef U Bachelor Chef U Bachelor Chef U
7 27 Bachelor Labor U Bachelor Labor U Bachelor Labor U Bachelor Labor U
12 40 Bachelor Teacher U Bachelor Teacher U Bachelor Teacher U * * *
13 37 Master Student U PhD Teacher U PhD Unemployed U PhD Unemployed U
14 39 PhD Teacher U PhD Teacher U PhD Teacher U PhD Unemployed U
16 24 Bachelor Waiter U Bachelor Waiter U Bachelor Cleaning U * * *
17 24 High school Deliveryman R High school U * * * * * *
@@
Ethics. The study protocol was reviewed and approved by the Institutional Review Board (IRB) of an anonymized university. All participants were fully informed of the research objectives, procedures, and potential impacts. They were reminded of their rights to withdraw at any time without justification, to request deletion of their data, and to review any information collected about them. All data were anonymized, and pseudonymous participant IDs were assigned. Interviewers underwent extensive training in qualitative methods and LGBTQ+ cultural competence to ensure respectful engagement. Each participant received 30 RMB (approximately \$4.50 USD) for each completed interview session. To better illustrate the longitudinal continuity and attrition in participants’ GSN app usage across the four data collection waves, Table 2 summarizes which gay social applications each participant used in 2014, 2017, 2020, and 2023.
Table 2. A summary of social applications used by the same cohort of participants used by participants across four years (2014, 2017, 2020, and 2023). The symbol ‘—’ indicates that missing data.
Participant 2014 2017 2020 2023
1 Aloha, Blued, Finka Ginder
2 Aloha, Blued, Finka Aloha, Blued Blued, Finka Wechat, Tinder
3 Aloha, Blued, Finka Blued
4 Aloha, Finka Aloha, Finka Blued, Aloha Hinge, Blued
5 Aloha, Blued, Finka Aloha, Blued Blued Hinge
6 Aloha, Blued, Finka Aloha, Blued, Finak Finka QSmatch
7 Aloha, Blued, Finka Aloha, Blued Blued, Finka QSmatch
8 Aloha, Finka Aloha, Finka Blued Blued
9 Aloha, Blued, Finka Blued Aloha, Finka Finka
10 Aloha, Finka Aloha, Finka Aloha, Finka QSmatch
11 Aloha, Blued, Finka Blued, Zank Aloha, Blued Blued
12 Aloha, Blued, Finka Blued, Finka Aloha, Finka
13 Aloha, Blued Aloha, Blued Aloha, Finka Blued
14 Aloha, Blued Aloha, Blued Aloha, Finka Blued
15 Blued, Zank Blued, Zank Aloha, Finka, Tinder Tinder, QSmatch
16 Blued Blued, Finka Blued, Aloha
17 Aloha, Finka Aloha, Finka
@@ 3.2. Procedure
We conducted four rounds of data collection in April 2014, 2017, 2020, and 2023. Each round consisted of two components: Step1: Participants complete questionnaires to provide a baseline assessment of their perceived usability of gay male dating apps. Step2: Participants take part in a qualitative, semi-structured interview regarding their attitudes toward gay dating apps, lasting approximately 45–60 minutes.
@@ 3.4. Data Analysis
Quantitative data were analyzed using descriptive statistics and exploratory factor analysis (EFA) in SPSS to identify attitude clusters. Survey results across the four time points were compared to examine longitudinal shifts. Qualitative data were transcribed and analyzed using thematic analysis, following Braun and Clarke’s six-phase framework. Coding followed an iterative process, our qualitative analysis was guided by the contextual integrity framework (grimm2010social; nissenbaum2004privacy; nissenbaum2009privacy). According to the framework, starting with open coding and then grouping by themes. Two trained researchers independently conducted line-by-line coding in NVivo 12 using a hybrid deductive–inductive strategy, where initial codes were informed by the research aims and latent codes emerged from the data. Inter-rater reliability was assessed on 20% of transcripts using Cohen’s Kappa ($\kappa=0.86$), indicating strong agreement, and discrepancies were resolved through consensus discussions. Data saturation was monitored in each wave and typically reached after 10–12 interviews. To track thematic evolution diachronically, we calculated code frequencies for each theme across 2014, 2017, 2020, and 2023. Pearson’s Chi-square tests ($\chi^{2}$) were used to determine the significance of thematic shifts, supporting the triangulation of qualitative results with the quantitative findings.
@@ 4.1. Quantitative Results
4.2. Qualitative Results
The qualitative findings largely corroborate the quantitative results. Initially, participants viewed gay social networking apps (GSN apps) as revolutionary tools for community building and identity affirmation. As a result, the apps were warmly received (2014–2017). From 2017 to 2020, technical architectures continued to improve. Basic usability also improved. Yet participants increasingly expressed disapproval on multiple fronts: commercialization; “fast-food” interactions; hookup culture; lookism; role-based discrimination; and pervasive deceptive practices. Reliance on GSN apps began to decline during this period. In the 2020–2023 interviews, disappointment became salient. Participants displayed clearer segmentation and a pragmatic stance. Many adopted international GSN apps (e.g., Grindr, Tinder). Use of domestic mainstream apps (e.g., Blued, Aloha) decreased. To ensure conceptual continuity across all four waves, we maintained a stable core codebook grounded in the research questions and prior literature. Two researchers re-coded a shared subset of earlier transcripts at each new wave to recalibrate interpretations of key concepts and prevent code drift. Any emergent refinements were applied retroactively to previous waves, ensuring that observed changes were due to genuine shifts in user experience rather than evolving coding criteria. Accordingly, $\chi^{2}$ statistics are used purely as descriptive indicators of theme prevalence under a consistent conceptual definition.